OpenStudy (anonymous):

show that f(x,y)=x^3*y is everywhere differentiable

OpenStudy (turingtest):

\[{\partial f \over \partial x}=3x^2y\]\[{\partial f \over \partial y}=x^3\]both x and y have domain \[(-\infty,\infty)\]in both derivatives, therefor f is differentiable everywhere.